OpenStudy (anonymous):

Can someone please help me solve my problem? I have a time limit I need to have this completed.

OpenStudy (anonymous):

An investor invested a total of $1,500 in two mutual funds. One fund earned a 6% profit while the other earned a 3% profit. If the investor's total profit was $51, how much was invested in each mutual fund?

OpenStudy (dumbcow):

A + B = 1500 .06A +.03B = 51 solve system for A,B i recommend multiplying 2nd equation by 100 to get rid of decimals

OpenStudy (anonymous):

@dumbcow okay, I can do that. I just wasn't sure how to set it up.

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

Thank you! :]

OpenStudy (anonymous):

I want to make sure that I am doing it correctly. when I multiply the second equation by 100 I will end up with 7a + 3b = 5100, right?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

oops i meant to put 6a + 3b = 5100

OpenStudy (dumbcow):

yes thats right